OpenStudy (anonymous):

I need help in figuring out the inequality formula for the following problem r/2 - 6 =14 Can anyone help me?

OpenStudy (liamschumm):

Hello there, Tiger!

OpenStudy (anonymous):

Hi can you help me

OpenStudy (liamschumm):

Try adding 6 to both sides.

OpenStudy (liamschumm):

Yep, that is what I am here for!

OpenStudy (liamschumm):

Ok, show me what you got.

OpenStudy (anonymous):

r/2 - 6 = 14 (6) r/2 -6 +14 = 20

OpenStudy (liamschumm):

Whoa, hold on.

OpenStudy (liamschumm):

You did well on the right side.

OpenStudy (liamschumm):

But you should be getting r/2 -6 + 6

OpenStudy (liamschumm):

r/2

OpenStudy (liamschumm):

So r/2 = 20

OpenStudy (anonymous):

ok thanks

OpenStudy (liamschumm):

So what does r=?

OpenStudy (anonymous):

40?

OpenStudy (liamschumm):

Yep, good job!

OpenStudy (liamschumm):

Thanks for the medal! Have a great day!

OpenStudy (anonymous):

Thank you so much